### v2

@@ 5.5 Nash equilibria in shortest-path games
We now consider shortest-path games. We show that from any NE, we can derive a finite-memory NE from the same initial vertex in which the same targets are visited and the cost profile of the derived NE is preferable to that of the original one. We construct NEs by extending the partially-defined Mealy machines described in Section 5.3.2 for outcomes given by Lemma 5.3 in another way than Section 5.4. We obtain NEs with strategies of memory size at most $n^{2}+2n$. We now fix a shortest-path game $\mathcal{G}=(\mathcal{A},(\mathsf{SPath}^{T_{i}}_{w})_{i\in\llbracket{}n\rrbracket{}})$ for the remainder of the section.
@@
Let $\sigma^{\prime}$ be an NE from a vertex $v_{0}$. There exists a finite-memory NE $\sigma$ from $v_{0}$ such that $\mathsf{SatPl}_{\mathcal{G}}(\mathsf{Out}(\sigma,v_{0}))=\mathsf{SatPl}_{\mathcal{G}}(\mathsf{Out}(\sigma^{\prime},v_{0}))$ and, for all $i\in\llbracket{}n\rrbracket{}$, $\mathsf{SPath}^{T_{i}}_{w}(\mathsf{Out}(\sigma,v_{0}))\leq\mathsf{SPath}^{T_{i}}_{w}(\mathsf{Out}(\sigma^{\prime},v_{0}))$ where each strategy of $\sigma$ has a memory size of at most $n^{2}+2n$. More precisely, we can bound their memory size by
$n\cdot(|\mathsf{VisPos}_{\mathcal{G}}(\mathsf{Out}(\sigma^{\prime},v_{0}))\setminus\{0\}|+2).$
Proof 5.14.
@@
We extend the Mealy machine of Section 5.3.2. We work with $I=\llbracket{}n\rrbracket{}$ in the following and drop $I$ from the notation to lighten it. Let $i\in\llbracket{}n\rrbracket{}$. We consider the Mealy machine $\mathcal{M}_{i}=(M,m^{\mathcal{D}}_{\mathsf{init}},\mathsf{up},\mathsf{nxt}_{i})$. We let $M=M^{\mathcal{D}}\cup\{\mathcal{P}_{i}\mid i\in\llbracket{}n\rrbracket{}\}$. The functions $\mathsf{up}$ and $\mathsf{nxt}_{i}$ extend $\mathsf{up}^{\mathcal{D}}$ and $\mathsf{nxt}_{i}^{\mathcal{D}}$ as follows. For all $(\mathcal{P}_{i^{\prime}},j)\in M^{\mathcal{D}}$ and $v\in V$ that does not occur in $\mathsf{sg}_{j}$, we let $\mathsf{up}((\mathcal{P}_{i^{\prime}},j),v)=\mathcal{P}_{i^{\prime}}$ and, if $v\in V_{i}$, we let $\mathsf{nxt}_{i}((\mathcal{P}_{i^{\prime}},j),v)=\tau_{-i^{\prime}}(v)$ if $i^{\prime}\neq i$ and $\mathsf{nxt}_{i}((\mathcal{P}_{i},j),v)$ is left arbitrary otherwise. For all $i^{\prime}\in\llbracket{}n\rrbracket{}$ and $v\in V$, we let $\mathsf{up}(\mathcal{P}_{i^{\prime}},v)=\mathcal{P}_{i^{\prime}}$ and, if $v\in V_{i}$, we let $\mathsf{nxt}_{i}(\mathcal{P}_{i^{\prime}},v)=\tau_{-i^{\prime}}(v)$ if $i^{\prime}\neq i$ and $\mathsf{nxt}_{i}(\mathcal{P}_{i},v)$ is left arbitrary.
We let $\sigma_{i}$ denote the strategy induced by $\mathcal{M}_{i}$. We have $|M|=n\cdot(k+2)$, therefore the memory size of $\sigma_{i}$ satisfies the announced upper bound. Furthermore, it follows from Lemma 5.8 that the outcome of $\sigma=(\sigma_{i})_{i\in\llbracket{}n\rrbracket{}}$ from $v_{0}$ is $\pi$ and $\sigma_{i}$ is based on $\mathcal{D}$ for all $i\in\llbracket{}n\rrbracket{}$.
@@
In this case, Theorem 5.13 provides the memory bound $2n$ if no players visit their target. However, the construction of Theorem 4.5 applies to such NEs in shortest-path games. We obtain therefore the more precise result.
Corollary 5.15.
If there exists an NE from $v_{0}$ such that no players see their target in its outcome, then there is a finite-memory NE from $v_{0}$ such that no players see their target in its outcome such that all strategies have a memory size of at most $n$.
@@ 6 Finite-memory Nash equilibria in safety games
We now move on to safety games. We show that from any NE, we can derive another finite-memory NE from the same initial vertex such that any safety objective satisfied by the outcome of the first NE also is satisfied by the outcome of the second one. In contrast to reachability games, we do not necessarily preserve the set of winning players and may obtain a superset thereof. We use the same construction as for shortest-path games and obtain the same arena-independent memory upper bounds.
@@ 6.1 Examples
Figure 6: Two-player arenas. Circle and squares are resp. $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ vertices.
@@
We consider the safety game played on the arena of Figure 6(a) such that the objectives of $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ are respectively $\mathsf{Safe}(\{v_{1},v_{4}\})$ and $\mathsf{Safe}(\{v_{5}\})$. The play $\pi=v_{0}v_{1}v_{2}v_{3}v_{4}^{\omega}$ is an NE outcome that only satisfies the objective of $\mathcal{P}_{2}$. As this play is a simple lasso, it admits a trivial simple segment decomposition.
@@
In reachability and shortest-path games, players have no incentive to stay within a segment of a play decomposition. This is no longer the case for safety games: stalling the progress of the play in a segment by exploiting the decomposition-based behaviour of the other players can allow a player to win by preventing a visit to undesirable vertices. We illustrate this property through the following example.
@@ 6.2 Well-structured Nash equilibrium outcomes
In this section, we formulate a variant of Lemma 5.3 for safety games. We transform an NE outcome into another NE outcome that is a simple play or a simple lasso. Let $\pi^{\prime}=v_{0}v_{1}\ldots$ be an NE outcome in $\mathcal{G}$. If it is possible to move from some vertex $v_{r}$ to a vertex $v_{\ell}$ with $\ell\leq r$, we let $\pi$ be a simple lasso of the form $v_{0}v_{1}\ldots(v_{\ell}\ldots v_{r})^{\omega}$ that exploits the earliest such available edge along $\pi^{\prime}$. Otherwise, we define $\pi=\pi^{\prime}$, as the absence of such edges implies that $\pi^{\prime}$ is already a simple play. The resulting play $\pi$ is an NE outcome by Theorem 4.3: target sets only occur in the common prefix of $\pi$ and $\pi^{\prime}$. By construction, there may be fewer vertices occurring in $\pi$ than in $\pi^{\prime}$ and thus $\mathsf{SatPl}_{\mathcal{G}}(\pi^{\prime})\subseteq\mathsf{SatPl}_{\mathcal{G}}(\pi)$ While this construction may enlarge the set of winning players, Example 6.2 illustrates that decomposition-based strategies are not sufficient in general to preserve a set of winning players exactly. To construct a decomposition-based NEs from the simplified play $\pi$, we will also require a well-chosen decomposition of $\pi$ (see Example 6.1). We use a decomposition analogous to those used in Lemma 5.3: we decompose $\pi$ into segments between two consecutive visits to targets.
@@
    if $k>0$, then for all $j\in\llbracket{}k\rrbracket{}$, there are no (simple) cycles made up of vertices other than $\mathsf{last}(\mathsf{sg}_{j})$ that appear in $\mathsf{sg}_{1}\cdot\ldots\cdot\mathsf{sg}_{j}$; and
@@
First, we assume that there exists some indices $\ell\leq r$ such that $(v_{r},v_{\ell})\in E$. We let $r^{\star}$ be the smallest index for which there exists such an $\ell$, and let $\ell^{\star}$ be the smallest among these. We define $\pi=v_{0}v_{1}\ldots(v_{\ell^{\star}}\ldots v_{r^{\star}})^{\omega}$, i.e., we take the first edge from a vertex in $\pi^{\prime}$ that leads to a previously visited vertex and loop in the resulting simple cycle. The play $\pi$ is a simple lasso and thus satisfies i. Its construction implies that iv also holds. By the characterisation of Theorem 4.3, $\pi$ is an NE outcome (as this characterisation applies to $\pi^{\prime}$).
We now define $k$ and the required simple decomposition $(\mathsf{sg}_{1},\ldots,\mathsf{sg}_{k+1})$ of $\pi$. By construction, we have $\mathsf{VisPos}_{\mathcal{G}}(\pi)\subseteq\mathsf{VisPos}_{\mathcal{G}}(\pi^{\prime})$. We let $k=|\mathsf{VisPos}_{\mathcal{G}}(\pi)\setminus\{0\}|\leq|\mathsf{VisPos}_{\mathcal{G}}(\pi^{\prime})\setminus\{0\}|$. If $k=0$, we consider the trivial decomposition $(\pi)$ of $\pi$. In this case, properties ii and iii hold trivially. We now assume that $k>0$ and write $\mathsf{VisPos}_{\mathcal{G}}(\pi)\cup\{0\}=\{\ell_{0}<\ell_{1}<\ldots<\ell_{k}\}$. For all $j\in\llbracket{}k\rrbracket{}$, we define $\mathsf{sg}_{j}=v_{\ell_{j-1}}\ldots v_{\ell_{j}}$ and we let $\mathsf{sg}_{k+1}=\pi{\ell_{k}}$. Property ii follows by definition of $(\mathsf{sg}_{1},\ldots,\mathsf{sg}_{k+1})$ and iii follows from the choice of $r^{\star}$: if there were a cycle witnessing the violation of iii, then the choice of $r^{\star}$ would no longer be minimal. This closes the first case.
@@ 6.3 Nash equilibria in safety games
We build on the NE outcomes given by Lemma 6.4 to show that we can derive, from any NE, another NE from the same initial vertex induced by Mealy machines of size quadratic in the number of players and whose outcome satisfies all of the safety objectives satisfied by the outcome of the original NE. We adapt the construction used for shortest-path games in the proof of Theorem 5.13: the only difference is that we need only monitor players whose objectives are not satisfied (as in reachability games), as the others cannot have a profitable deviation. To prove that the strategy profile we construct is an NE, we rely on property iii of outcomes obtained by Lemma 6.4. Intuitively, this property guarantees that players cannot halt the progress of the play in a segment of the decomposition on which we base the strategy to obtain a profitable deviation. Furthermore, players cannot have profitable deviations by leaving the current segment of the play: the other players switch to a punishing strategy of the player and, if the safety objective of the deviating player has not been falsified yet, then the NE outcome characterisation of Theorem 4.3 guarantees the effectiveness of the punishment by the other players.
@@
Let $\sigma^{\prime}$ be an NE from a vertex $v_{0}$ in $\mathcal{G}$. There exists a finite-memory NE $\sigma$ from $v_{0}$ such that $\mathsf{SatPl}_{\mathcal{G}}(\mathsf{Out}(\sigma^{\prime},v_{0}))\subseteq\mathsf{SatPl}_{\mathcal{G}}(\mathsf{Out}(\sigma,v_{0}))$ where each strategy of $\sigma$ has a memory size of at most $n^{2}+2n$. More precisely, we can bound their memory size by
$\max\{1,n-|\mathsf{SatPl}_{\mathcal{G}}(\mathsf{Out}(\sigma^{\prime},v_{0}))|\}\cdot(|\mathsf{VisPos}_{\mathcal{G}}(\mathsf{Out}(\sigma^{\prime},v_{0}))\setminus\{0\}|+2).$
@@
We extend the Mealy machine defined in Section 5.3.2. We let $I=\llbracket{}n\rrbracket{}\setminus\mathsf{SatPl}_{\mathcal{G}}(\pi)$ if this set is non-empty, and $I=\{1\}$ otherwise. Let $i\in\llbracket{}n\rrbracket{}$. We define a Mealy machine $\mathcal{M}_{i}=(M,m^{I,\mathcal{D}}_{\mathsf{init}},\mathsf{up},\mathsf{nxt}_{i})$ as follows. We let $M=M^{I,\mathcal{D}}\cup\{\mathcal{P}_{i}\mid i\in I\}$. The functions $\mathsf{up}$ and $\mathsf{nxt}_{i}$ extend $\mathsf{up}^{I,\mathcal{D}}$ and $\mathsf{nxt}_{i}^{I,\mathcal{D}}$ as follows. For all $(\mathcal{P}_{i^{\prime}},j)\in M^{I,\mathcal{D}}$ and $v\in V$ that does not occur in $\mathsf{sg}_{j}$, we let $\mathsf{up}((\mathcal{P}_{i^{\prime}},j),v)=\mathcal{P}_{i^{\prime}}$ and, if $v\in V_{i}$, we let $\mathsf{nxt}_{i}((\mathcal{P}_{i^{\prime}},j),v)=\tau_{-i^{\prime}}(v)$ if $i^{\prime}\neq i$ and $\mathsf{nxt}_{i}((\mathcal{P}_{i},j),v)$ is left arbitrary otherwise. For all $i^{\prime}\in\llbracket{}n\rrbracket{}$ and $v\in V$, we let $\mathsf{up}(\mathcal{P}_{i^{\prime}},v)=\mathcal{P}_{i^{\prime}}$ and, if $v\in V_{i}$, we let $\mathsf{nxt}_{i}(\mathcal{P}_{i^{\prime}},v)=\tau_{-i^{\prime}}(v)$ if $i^{\prime}\neq i$ and $\mathsf{nxt}_{i}(\mathcal{P}_{i},v)$ is left arbitrary otherwise.

### v3

@@ 5.5 Nash equilibria in shortest-path games
We now consider shortest-path games. We show that from any NE, we can derive a finite-memory NE from the same initial vertex in which the same targets are visited and the cost profile of the derived NE is preferable to that of the original one. We construct NEs by instantiating Mealy machines from Definition 5.13 for outcomes and decompositions given by Lemma 5.3 in another way than Section 5.4. We obtain NEs with strategies of memory size at most $n^{2}+2n$. We now fix a shortest-path game $\mathcal{G}=(\mathcal{A},(\mathsf{SPath}^{T_{i}}_{w})_{i\in\llbracket{}n\rrbracket{}})$ for the remainder of the section.
@@
Let $\sigma^{\prime}$ be an NE from a vertex $v_{0}$. There exists a finite-memory NE $\sigma$ from $v_{0}$ such that $\mathsf{SatPl}_{\mathcal{G}}(\mathsf{Out}(\sigma,v_{0}))=\mathsf{SatPl}_{\mathcal{G}}(\mathsf{Out}(\sigma^{\prime},v_{0}))$ and, for all $i\in\llbracket{}n\rrbracket{}$, $\mathsf{SPath}^{T_{i}}_{w}(\mathsf{Out}(\sigma,v_{0}))\leq\mathsf{SPath}^{T_{i}}_{w}(\mathsf{Out}(\sigma^{\prime},v_{0}))$ where each strategy of $\sigma$ has a memory size of at most
$n\cdot(|\mathsf{VisPos}_{\mathcal{G}}(\mathsf{Out}(\sigma^{\prime},v_{0}))\setminus\{0\}|+2)\leq n^{2}+2n$
Proof 5.23.
@@
We work with $I=\llbracket{}n\rrbracket{}$ in the following and drop $I$ from the notation of Definition 5.13 to lighten it. Let $i\in\llbracket{}n\rrbracket{}$. We define an $(\llbracket{}n\rrbracket{},\mathcal{D})$-compatible Mealy machine $\mathcal{M}_{i}=(M,(\mathcal{P}_{1},1),\mathsf{up},\mathsf{nxt}_{i})$ for $\mathcal{P}_{i}$ as follows. We let $M=M^{\mathcal{D}}\cup\{\mathcal{P}_{i}\mid i\in\llbracket{}n\rrbracket{}\}$. The functions $\mathsf{up}$ and $\mathsf{nxt}_{i}$ extend $\mathsf{up}^{\mathcal{D}}$ and $\mathsf{nxt}_{i}^{\mathcal{D}}$ as follows. For all $(\mathcal{P}_{i^{\prime}},j)\in M^{\mathcal{D}}$ and $v\in V$ that does not occur in $\mathsf{sg}_{j}$, we let $\mathsf{up}((\mathcal{P}_{i^{\prime}},j),v)=\mathcal{P}_{i^{\prime}}$ and, if $v\in V_{i}$, we let $\mathsf{nxt}_{i}((\mathcal{P}_{i^{\prime}},j),v)=\tau_{-i^{\prime}}(v)$ if $i^{\prime}\neq i$ and $\mathsf{nxt}_{i}((\mathcal{P}_{i},j),v)$ is left arbitrary otherwise. For all $i^{\prime}\in\llbracket{}n\rrbracket{}$ and $v\in V$, we let $\mathsf{up}(\mathcal{P}_{i^{\prime}},v)=\mathcal{P}_{i^{\prime}}$ and, if $v\in V_{i}$, we let $\mathsf{nxt}_{i}(\mathcal{P}_{i^{\prime}},v)=\tau_{-i^{\prime}}(v)$ if $i^{\prime}\neq i$ and $\mathsf{nxt}_{i}(\mathcal{P}_{i},v)$ is left arbitrary.
We let $\sigma_{i}$ denote the strategy induced by $\mathcal{M}_{i}$. We have $|M|=n\cdot(k+2)$, therefore the memory size of $\sigma_{i}$ satisfies the upper bound of the theorem. Furthermore, it follows from Lemma 5.14 that the outcome of $\sigma=(\sigma_{i})_{i\in\llbracket{}n\rrbracket{}}$ from $v_{0}$ is $\pi$ and $\sigma_{i}$ is based on $\mathcal{D}$ for all $i\in\llbracket{}n\rrbracket{}$.
@@ 6 Finite-memory Nash equilibria in safety games
We now move on to safety games. We show that from any NE, we can derive a finite-memory NE from the same initial vertex such that any safety objective satisfied by the outcome of the first NE also is satisfied by the outcome of the second one. In contrast to reachability games, we do not necessarily preserve the set of winning players and may obtain a superset thereof. We use the same construction as for shortest-path games and obtain the same arena-independent memory upper bounds.
@@ 6.1 Examples
Figure 6: Two-player arenas. Circle and squares are respectively $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ vertices.
@@
We consider the safety game played on the arena of Figure 6(a) such that the objectives of $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ are respectively $\mathsf{Safe}(\{v_{1},v_{4}\})$ and $\mathsf{Safe}(\{v_{5}\})$. The play $\pi=v_{0}v_{1}v_{2}v_{3}v_{4}^{\omega}$ is an NE outcome that satisfies only the objective of $\mathcal{P}_{2}$. As this play is a simple lasso, it admits a trivial simple segment decomposition.
@@
In reachability and shortest-path games, players have no incentive to stay within a segment of a play decomposition. This is no longer the case for safety games: stalling the progress of the play in a segment by exploiting the decomposition-based behaviour of the other players can allow a player to win by preventing a visit to undesirable vertices. We illustrate this property with the following example.
@@ 6.2 Well-structured Nash equilibrium outcomes
In this section, we formulate a variant of Lemma 5.3 for safety games. We transform an NE outcome into another NE outcome that is a simple play or a simple lasso. Let $\pi^{\prime}=v_{0}v_{1}\ldots$ be an NE outcome in $\mathcal{G}$. If it is possible to move from some vertex $v_{r}$ to a vertex $v_{\ell}$ with $\ell\leq r$, we let $\pi$ be a simple lasso of the form $v_{0}v_{1}\ldots(v_{\ell}\ldots v_{r})^{\omega}$ that exploits the earliest such available edge along $\pi^{\prime}$. Otherwise, we define $\pi=\pi^{\prime}$, as the absence of such edges implies that $\pi^{\prime}$ is already a simple play. The resulting play $\pi$ is an NE outcome by Theorem 4.3: target sets only occur in the common prefix of $\pi$ and $\pi^{\prime}$. By construction, there may be fewer vertices occurring in $\pi$ than in $\pi^{\prime}$ and thus $\mathsf{SatPl}_{\mathcal{G}}(\pi^{\prime})\subseteq\mathsf{SatPl}_{\mathcal{G}}(\pi)$ . While this construction may enlarge the set of winning players, Example 6.2 illustrates that decomposition-based strategies are not sufficient in general to preserve a set of winning players exactly. To construct a decomposition-based NEs from the simplified play $\pi$, we will also require a well-chosen decomposition of $\pi$ (see Example 6.1). We use a decomposition analogous to those used in Lemma 5.3: we decompose $\pi$ into segments between two consecutive visits to targets.
@@
    if $k>0$, then for all $j\in\llbracket{}k\rrbracket{}$, all cycles whose vertices all occur in $\mathsf{sg}_{1}\cdot\ldots\cdot\mathsf{sg}_{j}$ must contain $\mathsf{last}(\mathsf{sg}_{j})$; and
@@
First, we assume that there exist some indices $\ell\leq r$ such that $(v_{r},v_{\ell})\in E$. We let $r^{\star}$ be the smallest index for which there exists such an $\ell$, and let $\ell^{\star}$ be the smallest among these. We define $\pi=v_{0}v_{1}\ldots(v_{\ell^{\star}}\ldots v_{r^{\star}})^{\omega}$, i.e., we take the first edge from a vertex in $\pi^{\prime}$ that leads to a previously visited vertex and loop in the resulting simple cycle. The play $\pi$ is a simple lasso and thus satisfies i. Its construction implies that iv also holds. By the characterisation of Theorem 4.3, $\pi$ is an NE outcome (as this characterisation applies to $\pi^{\prime}$).
We now define $k$ and the required simple decomposition $(\mathsf{sg}_{1},\ldots,\mathsf{sg}_{k+1})$ of $\pi$. By construction, we have $\mathsf{VisPos}_{\mathcal{G}}(\pi)\subseteq\mathsf{VisPos}_{\mathcal{G}}(\pi^{\prime})$. We let $k=|\mathsf{VisPos}_{\mathcal{G}}(\pi)\setminus\{0\}|\leq|\mathsf{VisPos}_{\mathcal{G}}(\pi^{\prime})\setminus\{0\}|$. If $k=0$, we consider the trivial decomposition $(\pi)$ of $\pi$. In this case, properties ii and iii hold trivially. We now assume that $k>0$ and write $\mathsf{VisPos}_{\mathcal{G}}(\pi)\cup\{0\}=\{\ell_{0}<\ell_{1}<\ldots<\ell_{k}\}$. For all $j\in\llbracket{}k\rrbracket{}$, we define $\mathsf{sg}_{j}=v_{\ell_{j-1}}\ldots v_{\ell_{j}}$ and we let $\mathsf{sg}_{k+1}=\pi{\ell_{k}}$. Property ii follows by definition of $(\mathsf{sg}_{1},\ldots,\mathsf{sg}_{k+1})$ and iii follows from the choice of $r^{\star}$: if there were a cycle witnessing the violation of iii, then the choice of $r^{\star}$ would not have been minimal. This closes the first case.
@@ 6.3 Nash equilibria in safety games
We build on the NE outcomes given by Lemma 6.4 to show that we can derive, from any NE, another NE from the same initial vertex induced by Mealy machines of size quadratic in the number of players and whose outcome satisfies all of the safety objectives satisfied by the outcome of the original NE. We adapt the construction used for shortest-path games in the proof of Theorem 5.22: the only difference is that we need only monitor players whose objectives are not satisfied (as in reachability games), as the others cannot have a profitable deviation. To prove that the strategy profile that we construct is an NE, we rely on property iii of outcomes obtained by Lemma 6.4. Intuitively, this property guarantees that players cannot halt the progress of the play in a segment of the decomposition on which we base the strategy to obtain a profitable deviation. Furthermore, players cannot have profitable deviations by leaving the current segment of the play: the other players switch to a punishing strategy of the player and, if the safety objective of the deviating player has not been falsified yet, then the NE outcome characterisation of Theorem 4.3 guarantees the effectiveness of the punishment by the other players.
@@
Let $\sigma^{\prime}$ be an NE from a vertex $v_{0}$ in $\mathcal{G}$. There exists a finite-memory NE $\sigma$ from $v_{0}$ such that $\mathsf{SatPl}_{\mathcal{G}}(\mathsf{Out}(\sigma^{\prime},v_{0}))\subseteq\mathsf{SatPl}_{\mathcal{G}}(\mathsf{Out}(\sigma,v_{0}))$ where each strategy of $\sigma$ has a memory size of at most
$\max\{1,n-|\mathsf{SatPl}_{\mathcal{G}}(\mathsf{Out}(\sigma^{\prime},v_{0}))|\}\cdot(|\mathsf{VisPos}_{\mathcal{G}}(\mathsf{Out}(\sigma^{\prime},v_{0}))\setminus\{0\}|+2)\leq n^{2}+2n.$
@@
We let $I=\llbracket{}n\rrbracket{}\setminus\mathsf{SatPl}_{\mathcal{G}}(\pi)$ if this set is non-empty, and $I=\{1\}$ otherwise. Let $i\in\llbracket{}n\rrbracket{}$. We define an $(I,\mathcal{D})$-compatible Mealy machine $\mathcal{M}_{i}=(M,(\mathcal{P}_{\min I},1),\mathsf{up},\mathsf{nxt}_{i})$ for $\mathcal{P}_{i}$ as follows (cf. Definition 5.13). We let $M=M^{I,\mathcal{D}}\cup\{\mathcal{P}_{i}\mid i\in I\}$. The functions $\mathsf{up}$ and $\mathsf{nxt}_{i}$ extend $\mathsf{up}^{I,\mathcal{D}}$ and $\mathsf{nxt}_{i}^{I,\mathcal{D}}$ as follows. For all $(\mathcal{P}_{i^{\prime}},j)\in M^{I,\mathcal{D}}$ and $v\in V$ that does not occur in $\mathsf{sg}_{j}$, we let $\mathsf{up}((\mathcal{P}_{i^{\prime}},j),v)=\mathcal{P}_{i^{\prime}}$ and, if $v\in V_{i}$, we let $\mathsf{nxt}_{i}((\mathcal{P}_{i^{\prime}},j),v)=\tau_{-i^{\prime}}(v)$ if $i^{\prime}\neq i$ and $\mathsf{nxt}_{i}((\mathcal{P}_{i},j),v)$ is left arbitrary otherwise. For all $i^{\prime}\in\llbracket{}n\rrbracket{}$ and $v\in V$, we let $\mathsf{up}(\mathcal{P}_{i^{\prime}},v)=\mathcal{P}_{i^{\prime}}$ and, if $v\in V_{i}$, we let $\mathsf{nxt}_{i}(\mathcal{P}_{i^{\prime}},v)=\tau_{-i^{\prime}}(v)$ if $i^{\prime}\neq i$ and $\mathsf{nxt}_{i}(\mathcal{P}_{i},v)$ is left arbitrary otherwise.